Owens Company manufactures tennis racquets. The selling price per racquet averages $320 and variable costs per racquet are $190. The sales volume in dollars which produces a net income before taxes of $403,000 is $3,104,000. Calculate variable cost and fixed cost
Answer #1
Variable cost:
Sales volume in $=$ 3104000
selling price per racquet=$ 320
Sales volume in racquet=3104000/320=9700 racquets
variable costs per racquet=$ 190
Total variable cost=9700*190=$ 1843000
Fixed cost:
Cotribution=Sales volume in $-Total variable cost=3104000-1843000=$ 1261000
Contribution-Fixed cost=Net income
Fixed cost=Contribution-Net income=1261000-403000=$ 858000
I appreciate your ratings